OpenStudy (anonymous):

What is the slope-intercept form of the equation y – 7 = –5/2 (x + 4)?

OpenStudy (anonymous):

You need to manipulate the equation to be in the form \[y=mx+b\] 1) get y by itself 2) multiply the -5/2 through 3) simplify what you can.

OpenStudy (anonymous):

whatever freewilly922 said is very correct. I will show you the steps. \[y -5=-\frac{ 5 }{ 2 }x -10\]\[y =-\frac{ 5 }{2 }x -5\]This is the slope intercept form. Here,\[m =-\frac{ 5 }{ 2 }\]and\[b =-5\]